OpenStudy (anonymous):

Solve and show work h-48+17h=42 h=?

OpenStudy (anonymous):

ok so your goal is to get h by itself

OpenStudy (anonymous):

hmmm hmmmm

OpenStudy (amistre64):

combine like terms: h +17h = ? 42 +48 = ?

OpenStudy (anonymous):

combine like terms h+17h=18h your problem should look like this: 18h-48=42 add 48 to both sides 18h=90 divide each side by 18 h=5